OpenStudy (calculusxy):

MEDAL!!! Represent the area of a triangle with a base of 2x + y and a height of 4z. A. 2z(2x + y) B. 2x( z + y) C. 2xyz D. x(2x + y) E. 2z + 4xy

imqwerty (imqwerty):

area of triangle = 1/2 x base x height

OpenStudy (astrophysics):

Hint: \[A_{\triangle}=\frac{ 1 }{ 2 }bh\]

OpenStudy (astrophysics):

qwerty@!@@#@#

imqwerty (imqwerty):

:D lol

OpenStudy (astrophysics):

haha

OpenStudy (calculusxy):

I know that.

OpenStudy (calculusxy):

But I was thinking about substituting the variables for some numbers.

imqwerty (imqwerty):

base =2x+y height =4z

OpenStudy (astrophysics):

\[A = \frac{ 1 }{ 2 }(4z)(2x+y)\]

OpenStudy (astrophysics):

Simplify, all you had to do was plug it in

imqwerty (imqwerty):

A=2z(2x+y)

imqwerty (imqwerty):

do u wan the solution to that quadrilateral question @Astrophysics ??

OpenStudy (astrophysics):

YES

imqwerty (imqwerty):

ok

imqwerty (imqwerty):

lemme solve it XD

OpenStudy (calculusxy):

So why did it only cancel out the 4 and not the 2. was it to only find half of from only one part (or one number).

imqwerty (imqwerty):

ok can u tell what is 2+1??

imqwerty (imqwerty):

m not kiddin okay answer this^

OpenStudy (mrnood):

it is half base times height OR half height times base half height = 4z/2 = 2z

OpenStudy (phi):

it works the same as \[ \frac{1}{2} \cdot 4 \cdot 8 \] you can group it like this \[ \left(\frac{1}{2} \cdot 4 \right) \cdot 8\] only the 4 *1/2 is simplified to 2 and you get 2*8 or 16 if 8 were something complicated like (x+2y) it still be \[ \left(\frac{1}{2} \cdot 4 \right) \cdot (x+2y)\]

imqwerty (imqwerty):

we know that 2+1 = 3 nd if we find out (2+1)/2 then we can't cancel out 2 :) if u cancell it out then ur left with 1 but we know that 2+1 = 3 nd 3/2 is not 1. so u can't divide like that